OpenStudy (anonymous):

I will give a medal! Evaluate the function for x = –4c if c = –2. f(x)=-3x^2-4x

OpenStudy (ivettef365):

first substitute the c for -2, so what is -4(-2)

OpenStudy (nurali):

x = –4c c=-2 x=-4(-2) x=8 f(x)=-3x^2-4x f(8)=-3(8)^2-4(8) f(8)=-192-32 f(8)=-224